# Learning 3D Representations from Procedural 3D Programs

### v2

@@ Abstract
Self-supervised learning has emerged as a promising approach for acquiring transferable 3D representations from unlabeled 3D point clouds. Unlike 2D images, which are widely accessible, acquiring 3D assets requires specialized expertise or professional 3D scanning equipment, making it difficult to scale and raising copyright concerns. To address these challenges, we propose learning 3D representations from procedural 3D programs that automatically generate 3D shapes using simple primitives and augmentations.
Remarkably, despite lacking semantic content, the 3D representations learned from the procedurally generated 3D shapes perform on par with state-of-the-art representations learned from semantically recognizable 3D models (e.g., airplanes) across various downstream 3D tasks, including shape classification, part segmentation, and masked point cloud completion. We provide a detailed analysis on factors that make a good 3D procedural programs. Extensive experiments further suggest that current self-supervised learning methods on point clouds do not rely on semantics of 3D shapes, shedding light on the nature of 3D representations learned.
@@ 1 Introduction
Self-supervised learning (SSL) aims at learning representations from unlabeled data that can transfer effectively to various downstream tasks. Inspired by the success of SSL in language [10] and image representation learning  [14, 15], SSL for 3D point cloud understanding has gained considerable interest [37, 55, 26]. Recently, Point-MAE [26] and its follow-ups [59, 61, 41] exploit the masked autoencoding scheme for 3D point cloud representation learning, showing substantial improvements in various 3D shape understanding tasks (e.g., shape classification, part segmentation, and scene instance segmentation).
@@
However, unlike language and image data, which are abundantly available, 3D assets are less accessible due to the expertise required for creating 3D shapes using specialized software (e.g., Blender) or professional 3D scanners (e.g. LiDAR sensors). This scarcity of 3D shapes, known as data desert [11], limits the scalability of existing representation learning methods. Recent efforts have sought to expand 3D point cloud datasets at both the object level [8, 9] and scene level [3, 12, 53]. However, challenges unique to 3D data collection — such as copyright issues, format diversity, and scalability — remain unresolved.
@@
Figure 1: Self-supervised learning from (a) procedurally generated 3D shapes [45, 49, 50, 17] performs comparably to learning from (b) ShapeNet models that are semantically meaningful [4] across various downstream 3D understanding tasks. Both outperforms training from scratch significantly. In (c), the x-axis represents various tasks and benchmarks: ModelNet40 [22] and three variants of ScanObjectNN [34] for shape classification, and ShapeNetPart [54] for part segmentation.
To address these challenges, we explore learning point cloud representations from solely synthetic data generated via procedural 3D programs [45, 49, 50, 17] — examples are shown in Fig. 1a. Our data generation pipeline begins with sampling shapes from a set of simple 3D primitives (e.g., cubes, cylinders, and spheres). These primitives undergo affine transformations (e.g., scaling, translation, and rotation) and are combined to create diverse geometries. We then augment the composed shapes using predefined operations (e.g., Boolean operations) to further enhance topological diversity, and uniformly sample 3D surface points from these shapes for point cloud representation learning. We generate 150K synthetic 3D point clouds in  600 CPU hours, with a scalable pipeline capable of producing unlimited 3D shapes free of copyright concerns. Additionally, we generate 2K synthetic scenes, including room layouts and objects following the described steps, requiring  500 CPU hours. This leads us to two key questions:
  • Can existing self-supervised learning methods effectively capture rich and meaningful 3D representations from procedurally generated shapes alone?
  • How do these representations compare to those learned from human-crafted, semantically rich 3D models?
We investigate learning 3D representations exclusively from synthetic shapes generated via procedural programs, without relying on human-crafted data. To validate this approach, we conduct extensive benchmarks across representative self-supervised learning methods, including Point-MAE [26], Point-M2AE [59], PCP-MAE [61], and Masked Scene Contrast (MSC) [41]. Specifically, we evaluate 3D object SSL methods on tasks such as shape classification, part segmentation, and masked point cloud reconstruction, while 3D scene SSL methods are evaluated on semantic and instance segmentation tasks. Our main findings are:
  • Despite lacking semantics, SSLs from solely synthetic data achieve performance comparable to their ShapeNet-pretrained counterparts. We also validate the effectiveness of procedurally generated data in scene-level 3D understanding tasks (Fig. 1c, Tab. 1 - Tab. 5).
@@
  • We provide detailed insights into the factors that influence the quality of procedurally generated 3D datasets. We observe that learning performance improves notably with greater geometric diversity and increased dataset size (Tab. 4 and Fig. 5).
@@
  • Our in-depth analysis reveals structural similarities between 3D shape representations from models pretrained on synthetic 3D shapes and semantically meaningful 3D shapes, providing insights into the nature of the learned representations (t-SNE visualization in Fig. 7).
@@
To our best knowledge, this is the first systematic large-scale study on SSLs from procedural 3D programs. Our work is inspired by and builds upon a recent study that successfully trained large 3D reconstruction models exclusively on procedurally generated shapes [45, 17]. Our exploration is also closely related to prior efforts that learn image representations from procedural programs [2, 1]. Concurrent to our work, Yu et al. [56] demonstrated that procedurally generated videos can perform as effectively as natural videos for self-supervised video representation learning. Furthermore, our study is orthogonal yet complementary to recent efforts in scaling up 3D shape datasets [8, 9, 53].
@@ 2 Related Work
Figure 2: Learning from procedural 3D programs. (a) Synthetic 3D point clouds are generated by sampling, compositing, and augmenting simple primitives using procedural 3D programs [45]. (b) We experiment with multiple state-of-the-art self-supervised learning frameworks for learning 3D representations from synthetic data. Here, we illustrate the pretraining pipeline using Point-MAE [55], naming this variant Point-MAE-Zero, where “Zero” emphasizes the absence of any human-made 3D shapes. (c) We evaluate the pretrained models across various 3D shape understanding tasks.
3D Object Datasets. Large-scale datasets are essential in advancing 3D deep learning. Unlike 2D images or videos, building and annotating 3D models requires expertise with professional 3D software or scanning equipment, making the process more costly and time-intensive. Despite these challenges, significant efforts have been made to curate extensive 3D shape datasets [4, 43, 30, 32, 6, 12, 25, 39, 8]. For example, ShapeNet provides 3 million CAD models, with 51K of them being clean, high-quality models, which serves as the standard benchmark for training and evaluating models in the 3D object recognition community. A more significant limitation of ShapeNet is its strong bias towards rigid, man-made artifacts, reflecting the inherent bias of its source 3D model repositories. More recently, Objaverse-XL [9] expanded the 3D dataset to 10.2 million models, though scaling up has introduced challenges, including increased noise, format diversity, and unresolved copyright and legal issues. In contrast, we explore procedural 3D programs that generate 3D shapes from simple primitives. This synthetic approach can theoretically produce an unlimited number of 3D shapes without licensing concerns. Moreover, procedural 3D programs provide a principled approach to mitigating biases inherently present in manually curated datasets [1].
Supervised Learning for 3D Point Clouds. Unlike 2D images, 3D point clouds are inherently unordered and have an irregular structure, leading to extensive research on specialized neural network architectures [27, 28, 38, 62, 40, 42, 48, 19]. For instance, PointNet [27] introduces permutation-invariant operators and pooling layers feature aggregation across 3D points. PointNet++ [28] builds on this by incorporating a hierarchical spatial structure to capture more localized geometric information. DGCNN [38] adopts a graph-based approach, constructing a graph from input point clouds and applying a graph CNN to aggregate features from unordered 3D points. More recent models, such as the Point Transformer [62, 40, 42], utilize a modified transformer architecture tailored specifically for 3D point cloud processing. In our work, we employ a standard transformer architecture, aligning with recent trends in self-supervised learning [15, 26, 55].
Self-supervised Learning for Point Clouds. Self-supervised learning (SSL) for 3D representations aims to learn features that transfer well across diverse point cloud tasks, such as shape classification, object detection, and part segmentation. Recent SSL approaches for point clouds generally fall into two categories: contrastive learning [14, 46, 5] and masked autoencoding [37, 55, 26, 11, 29, 13, 60, 57, 58]. PointContrast [46] and DepthContrast [5] use an instance discrimination task [14] to learn 3D representations. OcCo [37] learns point cloud representations by reconstructing the original point cloud from occluded views, while IAE [51] trains an autoencoder to recover implicit features from point cloud inputs. Inspired by the success of masked autoencoding in vision [15] and language representation learning [10], Point-BERT [55] and Point-MAE [26] learn representations by predicting masked regions, using transformers as the underlying architecture. More works have been built upon Point-MAE: Point-M2AE [59] introduces a pyramid architecture for both the encoder and decoder to incorporate multi-scale and hierarchical information in 3D shapes. More recently, PCP-MAE [61] highlights the issue of centroid leakage, which trivializes the pretext task in the PointMAE framework, and proposes predicting centroids as an additional objective to strengthen masked point reconstruction. 3D scene SSLs directly takes in scene-centric point clouds is an exciting research direction [41, 16, 47, 36, 21]. In this work, we adopt PointMAE, Point-M2AE, PCP-MAE and MSC, as our primary self-supervised learning framework due to its state-of-the-art performance across various 3D understanding tasks. For ablation studies, we use Point-MAE as the baseline, as it serves as the foundational approach in this research direction and provides the most representative evaluation.
@@
Learning from Synthetic Data Synthetic data has become popular in computer vision, especially in scenarios where ground-truth annotations are difficult to obtain or where privacy and copyright issues arise. State-of-the-art performance in mid-level or 3D vision tasks is often achieved through training on synthetic data, including tasks like optical flow [33], depth estimation [52], dense tracking [18], relighting [49], novel view synthesis [50], and material estimation [20]. Procedurally generated synthetic data has also been explored for self-supervised representation learning in images [2, 1] and videos [56], and more recently for multi-view feed-forward 3D reconstruction [45]. Concurrent with our study, Yu et al. [56] investigates self-supervised video representation learning from procedurally generated images and videos. In this work, we explore self-supervised representation learning for point clouds, using synthetic 3D shapes generated by procedural 3D programs.
@@ 3 Learning from Procedural 3D Programs
We first introduce the procedural 3D programs [45, 49, 50] for generating unlimited number of synthetic 3D shapes using composition of simple primitive shapes (e.g., cylinders) and shape augmentation (Sec. 3.1). We then describe the masked autoencoding scheme [37, 55, 26] for learning 3D representations from synthetic 3D datasets (Sec. 3.2).
@@ 3.1 Procedural 3D programs
There is a line of work synthesizing procedural 3D shapes for vision tasks such as novel view synthesis [50], relighting [49], and material estimation [20]. Following recent methods [45] that use procedural 3D shapes for sparse-view reconstruction, we address self-supervised 3D representation learning from purely synthetic datasets. Fig. 2a illustrates our data pipeline:
@@
  • (2) Apply geometric augmentations (e.g., boolean differences, wireframe conversions) to enrich shape diversity (see [45] for details);
@@
In order to generate procedural 3D scenes, we follow MegaSynth [17] to procedurally generate 2K synthetic 3D scenes. Specifically, we first generate a floor plan and generate procedural 3D shapes with the above pipeline and place procedural 3D shapes in the scene based on the generated floor plan. We include details on how we generate 2K procedural 3D scenes in the supplementary material.
@@ 3.2 Procedural Pretraining
As depicted in Fig. 2b, we adopt Point-MAE [15], Point-M2AE [59], PCP-MAE [61] to train on procedurally generated 3D shapes. These methods rely on a masked autoencoding scheme [10, 15, 55], where the input point cloud is split into irregular patches and a large portion of them (60% by default) is randomly masked. A Transformer-based encoder–decoder network then attempts to reconstruct these masked patches, thereby learning 3D representations. Concretely, the reconstruction loss is computed as the $L_{2}$ Chamfer Distance between the predicted point patches $P_{\text{pre}}$ and the ground-truth patches $P_{\text{gt}}$:
$L=\frac{1}{|P_{\text{pre}}|}\sum_{a\in P_{\text{pre}}}\min_{b\in P_{\text{gt}}}\|a-b\|_{2}^{2}+\frac{1}{|P_{\text{gt}}|}\sum_{b\in P_{\text{gt}}}\min_{a\in P_{\text{pre}}}\|a-b\|_{2}^{2}.$ (1)
For scene-level SSLs, we adopt MSC [41], which combines masked auto-encoding and contrasive learning, to train on procedurally generated 3D scenes. While earlier self-supervised approaches [37, 55] often train on human-crafted 3D models (e.g., ShapeNet [4]), here we focus on purely synthetic data. We thus call add suffix -Zero to emphasize that it is trained exclusively on procedural 3D shapes with zero human-designed content. For clarity, we use the suffix -SN to indicate baselines trained on ShapeNet.
@@
We evaluate baselines on several 3D tasks, as summarized in Fig. 2c. For shape classification, we augment the pretrained Transformer encoder with a three-layer MLP classification head. For part segmentation, we aggregate features from the 4th, 8th, and final layers of the encoder, upsample them to all 2048 input points, and employ a segmentation head. For masked point cloud reconstruction, we use both the pretrained encoder and decoder with no architectural modifications. For scene-level methods, we use both instance segmentation and semantic segmentation finetuned from the pretrained SSLs with a linear prediction head. Detailed implementation settings are in the supplementary material.

### v3

# Semantic-Free Procedural 3D Shapes Are Surprisingly Good Teachers

@@ Abstract
Self-supervised learning has emerged as a promising approach for acquiring transferable 3D representations from unlabeled 3D point clouds. Unlike 2D images, which are widely accessible, acquiring 3D assets requires specialized expertise or professional 3D scanning equipment, making it difficult to scale and raising copyright concerns. To address these challenges, we propose learning 3D representations from procedural 3D programs that automatically generate 3D shapes using simple 3D primitives and augmentations.
Remarkably, despite lacking semantic content, the 3D representations learned from the procedurally generated 3D shapes perform on par with state-of-the-art representations learned from semantically recognizable 3D models (e.g., airplanes) across various downstream 3D tasks, such as shape classification, part segmentation, masked point cloud completion, and both scene semantic and instance segmentation. We provide a detailed analysis on factors that make a good 3D procedural programs. Extensive experiments further suggest that current 3D self-supervised learning methods on point clouds do not rely on semantics of 3D shapes, shedding light on the nature of 3D representations learned.
@@ 1 Introduction
Self-supervised learning (SSL) aims at learning representations from unlabeled data that can transfer effectively to various downstream tasks. Inspired by the success of SSL in language [devlin2018bert] and 2D images [he2020momentum, he2022masked], SSL for 3D point cloud understanding has gained considerable interest [wang2021occo, yu2022point, pang2022masked]. Recently, Point-MAE [pang2022masked] and its follow-ups [zhang2022point, zhang2024pcp, wu2023maskedscenecontrastscalable, wu2025sonata] exploit the masked autoencoding scheme [he2022masked] for 3D point cloud representation learning, showing substantial improvements in various 3D shape understanding tasks (e.g., shape classification, part segmentation, and scene instance segmentation).
@@
However, unlike language and image data, which are abundantly available on the Internet, 3D assets are less accessible, as their creation often requires domain expertise and specialized tools such as 3D modeling software (e.g., Blender) or scanning equipment (e.g. LiDAR sensors). This scarcity of 3D shapes, often referred to as the 3D data desert [dong2023autoencoderscrossmodalteacherspretrained], has significantly hindered the scalability of representation learning methods. Recent efforts have expanded point cloud datasets at both the object level [deitke2023objaverse, deitke2024objaverse] and the scene level [baruch2021arkitscenes, fu20213d, yeshwanth2023scannet++], but often rely on substantial human effort. Nevertheless, challenges unique to 3D data collection—such as copyright concerns, diverse file formats, and limited scalability—remain largely unresolved.
@@
Figure 1: Self-supervised learning from (a) procedurally generated 3D shapes [xie2024lrm, xu2018deep, xu2019deep, jiang2024megasynth] performs comparably to learning from (b) ShapeNet models that are semantically meaningful [chang2015shapenet] across various downstream 3D understanding tasks. Both outperforms training from scratch significantly. In (c), the x-axis represents various tasks and benchmarks: ModelNet40 [liu2019relation] and three variants of ScanObjectNN [uy2019revisiting] for shape classification, and ShapeNetPart [yi2016scalable] for part segmentation.
A common belief in 3D representation learning is that strong representations require semantically meaningful 3D shapes—objects such as chairs, airplanes, or indoor scenes—which are inherently costly to curate and difficult to scale. In this work, we challenge this assumption by asking a central question: Do we really need semantically meaningful 3D shapes to learn strong 3D representations?
To explore this, we investigate learning point cloud representations from purely synthetic data generated by procedural 3D programs [xie2024lrm, xu2018deep, xu2019deep, jiang2024megasynth], as illustrated in Fig. 1a. Our data generation pipeline begins by sampling simple 3D primitives (e.g., cubes, cylinders, spheres), which are then transformed via affine operations (e.g., scaling, translation, rotation) and composed into more complex geometries. We further apply augmentations such as Boolean operations to enrich topological diversity, followed by uniform surface point sampling to obtain point clouds suitable for representation learning. This approach is lightweight, efficient, and capable of generating unlimited number of 3D shapes with diverse geometric structures. Unlike standard datasets, our procedural shapes lack human-recognizable semantics.
We then pose a follow-up question: Are current 3D self-supervised learning (SSL) methods capable of leveraging such non-semantic, procedural shapes? To answer this, we generate 150K object-level and 4K scene-level synthetic 3D point clouds using our procedural programs, at a total cost of approximately 1400 CPU hours. Notably, the scale of our dataset exceeds that of widely used benchmarks such as ShapeNet (51K publicly available shapes) and ScanNet (1,513 indoor scenes). We benchmark multiple representative 3D SSL methods, including Point-MAE [pang2022masked] and its recent variants [zhang2022point, zhang2024pcp, wu2023maskedscenecontrastscalable], and evaluate them across a wide variety of downstream tasks such as shape classification, part segmentation, masked point cloud completion, and scene-level semantic and instance segmentation.
Our main findings are as follows:
  • Semantic-free procedural 3D shapes are surprisingly good teachers. Despite lacking semantic content, self-supervised models trained solely on synthetic data perform on par with counterparts trained on real 3D datasets such as ShapeNet [chang2015shapenet], Objaverse [deitke2023objaverse], and ScanNet [dai2017scannet]. Moreover, they significantly outperform models trained from scratch without any pretraining (Fig. 1c, Tabs. 15).
@@
  • Geometric diversity plays a key role in learning effective 3D representations. We provide detailed insights into the factors that influence the quality of procedurally generated 3D datasets. Our analysis shows that learning performance improves significantly with increased geometric diversity and larger dataset size (Tab. 4, Fig. 5).
@@
  • Current 3D SSL methods rely more on geometric cues rather than semantic content. Our in-depth analysis reveals strong structural similarities between representations learned from semantic-free synthetic procedural shapes and those learned from semantically meaningful 3D models (see t-SNE visualization in Fig. 7).
@@
To our best knowledge, this is the first systematic large-scale study on 3D SSLs from procedural 3D shapes. Our work is inspired by recent works that successfully train large 3D reconstruction models exclusively on procedurally generated shapes [xie2024lrm, jiang2024megasynth]. Our exploration is also closely related to prior efforts that learn image or video representations from procedural programs [baradad2021learning, baradad2022procedural, yu2024learning]. While recent efforts focus on scaling 3D datasets using human-designed models or 3D scans [deitke2023objaverse, deitke2024objaverse, yeshwanth2023scannet++], our approach is orthogonal and complementary, leveraging procedurally generated data to bypass manual design and scanning altogether.
@@ 2 Related Work
Figure 2: Learning from procedural 3D programs. (a) Synthetic 3D point clouds are generated by sampling, compositing, and augmenting simple primitives using procedural 3D programs [xie2024lrm]. (b) We experiment with multiple state-of-the-art self-supervised learning frameworks for learning 3D representations from synthetic data. Here, we illustrate the pretraining pipeline using Point-MAE [yu2022point], naming this variant Point-MAE-Zero, where “Zero” emphasizes the absence of any human-made 3D shapes. (c) We evaluate the pretrained models across various 3D shape understanding tasks.
3D Datasets. Significant efforts have been made to curate extensive 3D shape datasets [chang2015shapenet, modelnet40, reizenstein2021common, sun2018pix3d, collins2022abo, fu20213d, ma2024imagenet3d, wu2023omniobject3d, deitke2023objaverse]. For example, ShapeNet provides 3 million CAD models, with 51K high-quality shapes publicly available. More recently, Objaverse [deitke2023objaverse] and Objaverse-XL [deitke2024objaverse] expanded the 3D dataset to 10.2 million manually-created 3D shapes. However, challenges, such as format diversity and copyright and legal issues, remain unsolved. At the scene level, most commonly used datasets—ScanNet [dai2017scannet], Structured3D [structured3d], Matterport3D [matterport3d], nuScenes [nuscenes2019], SemanticKITTI [semantickitti], and Waymo [waymo]—are distributed under non-commercial or research-only licenses, limiting their applicability for broader use. A more recent effort, ASE [ASE], introduces 100,000 synthetic indoor scenes, also under a non-commercial license. In contrast, we explore procedural 3D programs that generate shapes and scenes from simple primitives, enabling the creation of unlimited synthetic objects and scenes without licensing constraints.
@@
Learning from Synthetic Data. Synthetic data has become popular in computer vision, especially in scenarios where ground-truth annotations are difficult to obtain or where privacy and copyright issues arise. State-of-the-art performance in mid-level or 3D vision tasks is often achieved through training on synthetic data, including tasks like optical flow [teed2020raft], depth estimation [yang2024depth], dense tracking [karaev2023cotracker], relighting [xu2018deep], novel view synthesis [xu2019deep], and material estimation [li2018learning]. Procedurally generated synthetic data has also been explored for self-supervised representation learning in images [baradad2021learning, baradad2022procedural] and videos [yu2024learning], and more recently for multi-view feed-forward 3D reconstruction [xie2024lrm]. In this work, we explore self-supervised representation learning for point clouds, using synthetic 3D shapes generated by procedural 3D programs.
Self-supervised Learning for Point Clouds. Recent self-supervised learning (SSL) methods for point clouds generally fall into two categories: contrastive learning [he2020momentum, xie2020pointcontrast, chhipa2022depth, gadelha2020label] and masked autoencoding [wang2021occo, yu2022point, pang2022masked, dong2023autoencoderscrossmodalteacherspretrained, qi2023recon, guo2023jointmae2d3djointmasked, zhang2022learning3drepresentations2d, zha2023compact3drepresentationspoint, zha2023towards]. Contrastive approaches such as PointContrast [xie2020pointcontrast] and DepthContrast [chhipa2022depth] rely on instance discrimination [he2020momentum] to learn view-invariant features. Inspired by the success of masked modeling in vision [he2022masked] and language [devlin2018bert], Point-BERT [yu2022point] and Point-MAE [pang2022masked] use transformer-based architectures to predict masked regions of the point cloud. Point-M2AE [zhang2022point] extends Point-MAE with a multi-scale pyramid design, while PCP-MAE [zhang2024pcp] addresses centroid leakage by adding centroid prediction as an auxiliary objective. Scene-level SSL methods that operate directly on large-scale point clouds have also gained traction [wu2023maskedscenecontrastscalable, hou2021exploringdataefficient3dscene, xie2020pointcontrastunsupervisedpretraining3d, Wang_2024_CVPR, liu2024pointcloudunsupervisedpretraining]. In this work, we adopt Point-MAE, Point-M2AE, PCP-MAE, and MSC as our primary SSL frameworks due to their strong performance on downstream 3D tasks. For ablation studies, we use Point-MAE as our baseline, as it represents a foundational and widely adopted approach in this line of research.
@@ 3 Learning from Procedural 3D Programs
We first introduce the procedural 3D programs [xie2024lrm, xu2018deep, xu2019deep] for generating unlimited number of synthetic 3D shapes using composition of simple primitive shapes (e.g., cylinders) and shape augmentation (Sec. 3.1). We then describe the masked autoencoding scheme [wang2021occo, yu2022point, pang2022masked] for learning 3D representations from synthetic 3D datasets (Sec. 3.2).
@@ 3.1 Procedural 3D programs
There is a line of work synthesizing procedural 3D shapes for vision tasks such as novel view synthesis [xu2019deep], relighting [xu2018deep], and material estimation [li2018learning]. Following recent methods [xie2024lrm] that use procedural 3D shapes for sparse-view reconstruction, we address self-supervised 3D representation learning from purely synthetic datasets. Fig. 2a illustrates our data pipeline:
@@
  • (2) Apply geometric augmentations (e.g., boolean differences, wireframe conversions) to enrich shape diversity (see [xie2024lrm] for details);
@@
In order to generate procedural 3D scenes, we follow MegaSynth [jiang2024megasynth] to procedurally generate 4K synthetic 3D scenes. Specifically, we first generate a floor plan and generate procedural 3D shapes with the above pipeline and place procedural 3D shapes in the scene based on the generated floor plan. We provide details on the generation of 4K procedural 3D scenes and visualizations of the generated shapes in the supplementary material.
@@ 3.2 Procedural Pretraining
We adopt Point-MAE [he2022masked], Point-M2AE [zhang2022point], PCP-MAE [zhang2024pcp] to train on procedural 3D shapes (Fig. 2b). These methods rely on a masked autoencoding scheme [devlin2018bert, he2022masked, yu2022point], where the input point cloud is split into irregular patches and a large portion of them (60% by default) is randomly masked. A Transformer-based encoder–decoder network then attempts to reconstruct these masked patches, thereby learning 3D representations. The reconstruction loss is computed as the $L_{2}$ Chamfer Distance between the predicted point patches $P_{\text{pre}}$ and the ground-truth patches $P_{\text{gt}}$:
$L=\sum_{x\in\{P_{\text{pre}},P_{\text{gt}}\}}\frac{1}{|x|}\sum_{a\in x}\min_{b\in x^{\prime}}\|a-b\|_{2}^{2}$ (1)
where $x^{\prime}=P_{\text{gt}}$ if $x=P_{\text{pre}}$, and vice versa. For scene-level SSLs, we adopt MSC [wu2023maskedscenecontrastscalable], which combines masked auto-encoding and contrasive learning, to train on procedurally generated 3D scenes.
@@
We evaluate baselines on several 3D tasks, as summarized in Fig. 2c. For shape classification, we augment the pretrained Transformer encoder with a three-layer MLP classification head. For part segmentation, we aggregate features from the 4th, 8th, and final layers of the encoder, upsample them to all 2048 input points, and employ a segmentation head. For masked point cloud reconstruction, we use both the pretrained encoder and decoder with no architectural modifications. For scene-level methods, we use both instance segmentation and semantic segmentation finetuned from the pretrained SSLs. Detailed implementation settings are in the supplementary material.